OpenStudy (lilai3):

Please help!!! An irregular heptagon has angles that measure x degrees, x degrees, 2x degrees, 2 x degrees, 3x degrees, 3x degrees, and 4x degrees. What is the measure of x?

OpenStudy (jennychan12):

heptagon = 7 sides interior angle formula : 180(n-2) n is number of sides so 6(180) 1080 15x = 1080 FIND X

OpenStudy (mathstudent55):

A heptagon is a 7-sided polygon. The sum of the measures of the interior angle of a ploygon is (n - 2)180, where n is the number of sides. x + x + 2x + 2x + 3x + 3x + 4x = 16x (7 - 2)180 = 5(180) = 900 16x = 900 Solve for x